# Research on Superalignment Should Advance Now with Parallel Optimization of Competence and Conformity

### v1

@@ 6 Conclusion
In this paper, we argue that superalignment is achievable and research on it should advance immediately, through simultaneous and alternating optimization of task competence and value conformity. We provide a formal definition of superalignment, originating from scalable oversight. Then, we review three existing paradigms and explore their limitations. To address the limitation, we propose two fundamental principles as a potential path to superalignment. First, determining an appropriate capability and capacity gap. Second, increasing the diversity of signals. We then outline a conceptual framework for superalignment, including a potential evaluation method to align ASI. By laying this foundation, we aim to illuminate the landscape for addressing the critical challenge of aligning future ASI with human values, ensuring its benefits for humanity.
Impact Statement
This work introduces a formal definition of superalignment, connecting existing concepts by exploring their interconnections and limitations. We then propose a potential pathway to achieve superalignment, grounded in two fundamental principles: (1) determining an appropriate capability and capacity gap and (2) increasing the diversity of signals.
Although our work proposes a potential pathway to address key challenges, it still contains three main challenges that should be addressed in the future. First, our position relies on assumptions about the speculative nature of ASI, which has not yet been realized. This introduces uncertainty regarding the applicability of our methods in future scenarios. Second, our evaluations are limited in terms of model size scalability. While we have demonstrated promising results that address current limitations and highlight the potential of our approach, our experiments are constrained by the use of a 7B parameter model as the largest tested model. Lastly, as shown in Fig. 1, it is important to consider not only $ASI_{unaligned}$ and $ASI_{aligned}$ but also cases where ASI is partially aligned with specific objectives—for example, an ASI model that is aligned with ethics but unaligned with safety. Active discussion is needed to deepen our understanding of ASI alignment.
This paper could influence organizations considering the development of superalignment AI models by not only clarifying existing alignment methods and their limitations but also presenting a potential pathway to achieving superalignment. By addressing current challenges and proposing a potential solution, we aim to contribute to the broader goal of developing AI models that provide benefits to humanity while minimizing potential harm.
@@ Appendix A Appendix
A.1 Potential Directions in Detail
@@
Encouraging data diversity can play a pivotal role in fostering creativity and robustness in alignment strategies. Multi-AI system collaboration, where AI systems with different specializations or perspectives interact, has shown promise in enabling emergent behaviors and problem-solving approaches beyond the capabilities of individual AI systems (Guo et al., 2024). Competition among agents can also drive improvement, as adversarial settings often reveal diverse argument data that finds alignment weaknesses and suggests novel pathways for improvement (Khan et al., 2024). However, diversity-based approaches require careful curation and filtering to ensure that resulting behaviors are beneficial and do not inadvertently reinforce undesirable outcomes.
@@
Iterative teacher-student training involves a cyclical process where a more advanced teacher guides a less capable student, and the teacher subsequently obtains feedback from the student for adjustments (Yuan et al., 2021). This process can be extended such that the student becomes the teacher in the next iteration, guiding the previous teacher to improve its alignment capabilities. This approach enables incremental learning, where each output generated by the AI system builds upon and enhances the alignment capabilities of its predecessor. While this method offers scalability and the potential to bootstrap stronger AI systems from weaker supervision signals, challenges such as those posed by W2SG (Sec. 3.3) and self-enhancement (Sec. 3.2) may arise due to limited diversity. Effective strategies must be devised to mitigate these risks.
@@
Search-based methods use optimization techniques to explore the space of possible alignment strategies, seeking an optimal pathway to systematically navigate complex alignment challenges (Khanov et al., 2024). These methods are particularly useful when the optimal alignment direction is unclear, as they allow for exploration and experimentation in uncertain spaces. Techniques like Monte Carlo tree search (undefinedwiechowski et al., 2022) or gradient-free optimization (Larson et al., 2019) can guide this process by prioritizing regions of the search space with the highest potential for success. However, these methods require significant computational resources.
A.2 Experiment Setup
A.2.1 General
@@
For sandwiching, we use the OpenAI API for the debate experiment. For self-enhancement and weak-to-strong generalization, we fully fine-tune all model variants using 16 NVIDIA A100 GPUs (80GB each), AMD EPYC 7V12 64-Core Processor CPU, with 1.73TB memory. To reduce memory redundancy during training, we utilize DeepSpeed ZeRO-3 (Rajbhandari et al., 2020) alongside Accelerate (Gugger et al., 2022), using bfloat16 half-precision. We use GPT-2 (Radford et al., 2019), Phi-3.5 (Abdin et al., 2024), Qwen2.5 (Yang et al., 2024a), and GPT-4 (OpenAI et al., 2024) series for our evaluation. For all the model information, please refer to Tab. A.4.
A.2.2 Sandwiching
@@
For the sandwiching experiment, we conduct a debate experiment to evaluate different capabilities in judge-debater and debater-debater setups. In the judge-debater setup, a previous study by Khan et al. (2024) demonstrated that the debater is more persuasive when advocating for the correct answer across various debate capabilities. Our goal is to evaluate the effectiveness of debates when the debater and judge have different levels of capability by directly comparing the judge’s accuracy and quality of debate statements.
@@
We use two proprietary models: GPT-4-Turbo and GPT-4o with API access. GPT-4o is known to be a higher-capability model compared to GPT-4-Turbo (OpenAI, 2024a). We test two different debaters arguing for their assigned answer. Where one debater argues for a correct answer (Helpful), and another debater argues for an incorrect answer (Sneaky). We run 47 randomly chosen GPQA (Rein et al., 2024) and QuALITY (Pang et al., 2022) binary QA converted test set for the evaluation with three debate turns. We run each setup four times and calculate the mean values. For QuALITY, we use the debater and judge prompts as described by Khan et al. (2024). For GPQA, we adopt a similar prompt style based on the QuALITY prompts. Detailed templates for GPQA prompts are provided in Tab. A.4, Tab. A.4, Tab. A.4, and Tab. A.4.
@@
A.2.3 Self-Enhancement
For self-enhancement setup, we utilize Learning from AI Feedback (LAIF) training with Direct Preference Optimization (DPO) (Rafailov et al., 2023), we adopt a self-training setup to evaluate the performance when there is no supervision from AI systems with greater capabilities. Instead, the supervision comes from models with equal capabilities.

### v2

@@ 6 Conclusion
In this paper, we argue that superalignment is achievable and research on it should advance immediately, through alternating optimization of competence and conformity. We provide its formal definition rooted in scalable oversight and the capability–capacity gap, and review existing paradigms to expose their limitations. To address these, we propose two novel principles. (1) Determining an appropriate gap, and (2) increasing the diversity of signals, and then outline the ACCO framework, with a potential evaluation method to align ASI. By laying this foundation, we aim to illuminate the landscape for addressing the critical challenge of aligning future ASI with human values.
@@ Appendix A Appendix
A.1 Alternative Views
Superalignment is Unattainable  Absence of ASI and the lack of informed oversight have led some to argue that superalignment is unattainable due to $C(\mathcal{A})\gg C(\mathcal{H})$ (Sec. 2). Given the current limitations, humans cannot reliably supervise AI systems $\mathcal{A}$ when their capabilities exceed human expertise, making it difficult to prevent $\mathcal{A}_{\text{ASI}}$ from deceiving its weaker teacher [31]. Moreover, while existing alignment methods [204, 148, 71, 196, 66] offer limited success in aligning $\mathcal{A}_{\text{ANI}}$, they face scalability and reliability issues. However, we show in Sec. 4.4 (Superalignment Evaluation) that $\mathbb{D}_{U}$ can be approximated without golden label.
Superalignment is an Emergent Property  Building on scaling laws [93, 186, 128], one could argue that superalignment might emerge as a byproduct of intelligence expansion rather than requiring proactive intervention. However, concerns about deceptive alignment [33, 75] and inverse scaling [117] persist. Formally, increased capacity $C(\mathcal{A})$ does not guarantee aligned capability $U(\mathcal{A})$, and superalignment explicitly requires minimizing the capability-capacity gap $|\mathbb{D}_{U}-\mathbb{D}_{C}|$. As models develop advanced capability, new challenges arise [174, 32, 68]. Relying solely on emergent properties poses risks, including unpredictability in AI behavior.
[…]
No Fundamental Difference from Current Alignment  Superalignment can simply be seen as advanced alignment, relying on current feedback techniques [74, 58]. We argue instead that superalignment arises when $C(A)\gg C(H)$ and humans can no longer supervise effectively. Unlike conventional alignment which minimize capability gap $|U(\mathcal{A})-U(\mathcal{H})|$, our framework addresses the divergence between capability-capacity $|\mathbb{D}_{U}[U(\mathcal{A}),U(\mathcal{H})-\mathbb{D}_{C}[C(\mathcal{A}),C(\mathcal{H})|$, requiring long-term, cross-generational methods beyond typical fine-tuning. For more details about the difference between superalignment and standard alignment, please refer to Appendix A.3.
Superalignment Can Be Achieved through Inference Scaling  Some argue that superalignment may emerge from inference-time scaling rather than model scaling [127, 201]. While inference scaling shows promise on structured tasks [69], and may improve performance $U(\mathcal{A})$, it does not increase model capacity or internalized knowledge $C(\mathcal{A})$, thus violating Premise 1. As a result, the capability–capacity gap $|\mathbb{D}_{U}-\mathbb{D}_{C}|$, under the assumption $C(\mathcal{A})\gg C(\mathcal{H})$, cannot be minimized.
Our Counterargument  As discussed in Sec. 2.2, we state our position that superalignment is achievable and research on it should advance immediately. Although current paradigms have limitations, delaying research would pose more significant risks [54, 79, 21, 67]. Instead, we propose a simultaneous optimization approach to achieve superalignment, integrating task competence [125] and value conformity [7, 174, 131].
A.2 Impact Statement
This work introduces a formal definition of superalignment, connecting existing concepts by exploring their interconnections and limitations. We then propose a potential pathway to achieve superalignment, grounded in two fundamental principles: (1) determining an appropriate capability-capacity gap and (2) increasing the diversity of signals.
Although our work proposes a potential pathway to address key challenges, it still contains three main challenges that should be addressed in the future. First, our position relies on assumptions about the speculative nature of ASI, which has not yet been realized. This introduces uncertainty regarding the applicability of our method in future scenarios, but we argue its relevance for work that approaches ASI by dominating human performance across multiple criteria and contexts. Second, our evaluations are limited in terms of model size and scalability. While we demonstrate promising results that address current limitations and highlight the potential of our approach, our experiments are constrained by the use of a 7B parameter model as the largest one tested. Lastly, as shown in Fig. 1, it is important to consider both aligned and unaligned $\mathcal{A}_{\text{ASI}}$ but also cases where ASI is partially aligned with specific objectives—for example, an ASI model aligned with ethics but unaligned with safety. Active discussion is needed to deepen the collective understanding of ASI alignment.
In terms of ethical concerns, (1) Amplified Biases. W2SG may propagate biases inherent in training data or weak models. (2) Increased Safety Risks. More capable models are more susceptible to noise, which may trigger unintended harmful behaviors. (3) Other Impacts AI’s biases and potential sycophancy can produce outputs overly aligned with unreliable human feedback, causing social harm.
We hope that this paper will influence organizations considering the development of superalignment AI models by not only clarifying existing alignment methods and their limitations but also presenting a potential pathway for achieving superalignment. By addressing current challenges and proposing a potential solution, we aim to contribute to the broader goal of developing AI models that provide benefits to humanity while minimizing potential harms.
A.3 Distinguishing Superalignment from Existing Alignment
In this section, we clarify the core differences between superalignment and existing alignment, especially in terms of their assumptions, goals, and methodological constraints.
Definition and Objective Existing alignment aims to align an AI system’s behavior with human values or preferences i.e., $U(\mathcal{A})\approx U(\mathcal{H})$. This assumes the model has sufficient capacity such that $C(\mathcal{A})\geq C(\mathcal{H})$ and a utility that is comparable to or lower than that of humans such that $U(\mathcal{A})<U(\mathcal{H})$. As formalized in Def. 2.1 in Sec. 2.1, the objective is to minimize the utility gap, defined by $\mathcal{A}^{*}=\mathop{\text{argmin}}\nolimits_{\mathcal{A}}|U(\mathcal{A})-U(\mathcal{H})|$. In contrast, superalignment aims to align an AI system whose utility $U(\mathcal{A})$ is required to significantly exceed that of any human, i.e., $U(\mathcal{A}_{\text{ASI}})\gg U(\mathcal{H})$. In such cases, human capability $U(\mathcal{H})$ is no longer a valid alignment reference. Instead, the superalignment objective focuses on minimizing the gap between a model’s measured capability and capacity, quantified by $\mathbb{D}_{U}$ and $\mathbb{D}_{C}$, respectively. As introduced in Def. 2.1, this is formalized as $\mathcal{A}^{*}=\mathop{\text{argmin}}\nolimits_{\mathcal{A}}\left|\mathbb{D}_{U}[U(\mathcal{A}),U(\mathcal{H})]-\mathbb{D}_{C}[C(\mathcal{A}),C(\mathcal{H})]\right|$ without changing $C(\mathcal{A})$.
Key Differences The core difference lies in the supervision approach. Existing alignment assumes that humans can provide or verify supervision signals, making techniques such as fine-tuning or reinforcement learning from human feedback (RLHF) feasible. Superalignment, however, arises in contexts where human supervision fails; e.g., when evaluating a model’s proof of the Riemann Hypothesis, even expert humans may be unable to assess the correctness of the solution. Consequently, existing alignment focuses on aligning models to human-level utility, $|U(\mathcal{A})-U(\mathcal{H})|<\epsilon$, while superalignment aims to align models to their own capacity in a way that conforms to intended utility structures, $|\mathbb{D}_{U}[U(\mathcal{A}),U(\mathcal{H})]-\mathbb{D}_{C}[C(\mathcal{A}),C(\mathcal{H})]|<\epsilon$. The two also differ in terms of risk. Existing alignment assumes that the human-estimated utility $\hat{U}_{\mathcal{H}}(\mathcal{A})$ is a reliable proxy for the golden utility $U(\mathcal{A})$, and that misalignment primarily results from human misjudgment or bias, with the deviation $|\hat{U}_{\mathcal{H}}(\mathcal{A})-U(\mathcal{A})|$ remaining sufficiently small. In contrast, superalignment assumes in domains where $\hat{U}_{\mathcal{H}}(\mathcal{A})$ is undefined or unreliable, such that $|\hat{U}_{\mathcal{H}}(\mathcal{A})-U(\mathcal{A})|$ becomes large due to emergent capabilities that are unpredictable or unverifiable by humans (e.g., treacherous turn).
Evaluation In terms of evaluation, while existing alignment assumes access to a ground truth $\mathbf{y}$, superalignment lacks such ground truth, which is often unavailable or unassessable, making its evaluation particularly challenging. To lay the groundwork for evaluating superalignment, we propose several approximation methods in Sec. 4.4.
A.4 Potential Directions in Detail
The conceptual framework described in Sec. 4.3 is promising, but challenges remain in implementing it in practice. In this section, we provide detailed potential directions.
@@ A.4 Potential Directions in Detail
Data diversity is essential to avoid mode collapse (Principle 2). Formally, let the supervision signal be denoted as $\mathcal{S}=\{\mathbf{x}_{i},\mathbf{y}_{i}^{j}\}$, where $\mathbf{y}_{i}^{j}\sim\mathcal{A}_{j}(\mathbf{x}_{i})$ for models $\mathcal{A}_{j}$ with varying capacities or perspectives, or $\mathbf{y}_{i}^{j}\sim\mathcal{H}_{j}(\mathbf{x}_{i})$ for human agents with diverse backgrounds. Diversity in $\mathcal{S}$ helps reduce the bias in $\hat{U}(\mathcal{A})=\mathbb{E}_{\mathbf{x}\sim\mathcal{X}}[\hat{U}(\mathcal{A}(\mathbf{x}))]$. Collaboration among heterogeneous AI systems [72] and competition among models [95] expose alignment weaknesses and solutions. However, filtering is essential to prevent undesirable behaviors and avoid noisy signals that may increase $|\mathbb{D}_{U}-\mathbb{D}_{C}|$, as discussed in Sec. 3.1 and Sec. 3.3.
@@
In iterative teacher-student training when $\mathcal{C}(\mathcal{A}_{i-1})<\mathcal{C}(\mathcal{A}_{i})$, a strong teacher $\mathcal{A}^{t-1}_{i}$ generates a solution $\mathbf{y}^{t-1}_{i}$ to supervise a weaker student $\mathcal{A}^{t-1}_{i-1}$, and the supervised weaker student $\mathcal{A}^{t}_{i-1}$ then provides feedback $\mathbf{y}^{t}_{i-1}$ to improve the teacher $\mathcal{A}^{t}_{i}$ [203]. This cyclical process can enhance alignment but may face challenges from limited diverse feedback (Sec. 3.3 and Sec. 3.2).
@@
Search-based methods optimize alignment strategies by exploring possible alignment strategies [96]. Given a fixed model capacity $C(\mathcal{A})$, these methods aim to optimize the realized capability $U(\mathcal{A})$ by searching over policy or alignment configurations to minimize the gap $|\mathbb{D}_{U}-\mathbb{D}_{C}|$. Techniques such as Monte Carlo tree search [180] or gradient-free optimization [104] can guide promising directions. However, these methods may require high computational resources.
A.5 Superalignment Evaluation in Detail
[…]
Extrapolation:  Uncovering emergent behaviors and potential misalignments as AI scales [117]. This method continuously evaluates the model as it scales to detect performance shifts. Perez et al. [141] propose using AI to generate evaluation datasets, identifying unexpected behaviors. Additionally, we suggest tracking alignment failures when scaling, as misalignments may follow a U-shaped pattern, emerging after initial peaks or valleys in tuning. Another extrapolation-based method may involve training AI models on easy tasks and evaluating their generalization to harder problems [173]. Formally, let $\{\mathcal{A}_{1},\dots,\mathcal{A}_{K}\}$ be $K$ models with increasing capacities, i.e., $C(\mathcal{A}_{1})<\cdots<C(\mathcal{A}_{K})$. Let $\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}$ be a corresponding sequence of tasks with increasing complexity. For each model-task pair $(\mathcal{A}_{i},\mathbf{x}_{i})$, the capability can be approximated as $\hat{U}(\mathbf{y}_{i})$. Extrapolation examines whether $\{\hat{U}(\mathbf{y}_{i})\}_{i=1}^{K}$ exhibits a consistent improvement or U-shaped pattern, revealing potential misalignment as models scale.
Peer Review:  Domain experts can qualitatively assess model outputs and reasoning for alignment. Peer review, shown to enhance academic performance [48], has also been used to mitigate noisy supervision [112]. Similarly, expert panels can evaluate a paper draft on technical accuracy and clarity. Engaging multiple reviewers ensures a more objective assessment of a model’s performance and alignment with human values. Formally, let $\{\mathcal{H}_{1},\dots,\mathcal{H}_{J}\}$ be $J$ reviewers with their own evaluation function $\hat{U}_{\mathcal{H}_{i}}$. For input $\mathbf{x}$, $\mathbf{y}=\mathcal{A}(\mathbf{x})$ is evaluated by all reviewer $\hat{U}(\mathbf{y})=\frac{1}{J}\sum_{i=1}^{J}\hat{U}_{\mathcal{H}_{i}}(\mathbf{y})$.
A.6 A Comprehensive Overview of Limitations in Existing Paradigms
Sandwiching
Despite promising results, Debate has fundamental limitations. (1) Reliance on Human Capability: Humans must be able to objectively evaluate AI-generated arguments, insights, or partial solutions and effectively handle task decomposition and integration when they cannot directly assess or provide the final solution. In short, human should reliably provide golden utility of AI-generated outputs $\hat{U}_{\mathcal{H}}(\mathbf{y})\approx U(\mathbf{y})$. However, for superhuman task $\mathbf{x}$ where $C(\mathcal{A})\gg C(\mathcal{H})$, the gap $|\hat{U}_{\mathcal{H}}(\mathbf{y})-U(\mathbf{y})|$ is large, making human supervision unreliable. See empirical evidence in Tab. 3.1, which shows that humans might be misled by AI debaters. (2) Task Requirements: The task must be understandable to humans (for identifying knowledge or task decomposition), decomposable (for recursive amplification), and at least partially solvable by AI (e.g., completing sub-tasks). However, this may be infeasible in the superalignment setting, where both the AI’s capacity far exceed human capacity, i.e., $C(\mathcal{A}_{\text{ASI}})>C(\mathcal{A})\gg C(\mathcal{H})$. (3) Supervision Signal Noise: Uncertainty in both humans and AI [176] means optimal solutions cannot always be elicited. Besides, biases in both humans and AI, along with AI’s sycophancy [163], can lead to over-alignment with unreliable human feedback or deception [75]—especially when humans fail to detect AI’s flaws. Concretely, the quality of the signal $\mathcal{S}=\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ deteriorates when both human and AI signals are noisy or biased, increasing capability-capacity gap $|\mathbb{D}_{U}-\mathbb{D}_{C}|$, and resulting in sycophancy or deception.
Self-Enhancement
Similarly, several limitations also lie in the Self-Enhancement paradigm. (1) Error Accumulation [50]: Because supervision signals $\mathcal{S}=\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ are entirely derived by AI, any errors or noise can propagate and accumulate over time, increasing capability-capacity gap $|\mathbb{D}_{U}-\mathbb{D}_{C}|$ potentially leading to issues like polarization [142, 102]. Tab. LABEL:tab:aif_result in Appendix shows that continual training on self-generated data progressively degrades performance as model size increases. (2) Self-Preference [185]: Because AI is inherently biased towards self-generated content, it results in suboptimal critique and verification performance (false positives). Formally, in an iterative refinement process where a model $\mathcal{A}_{i-1}$ critiques its successor $\mathcal{A}_{i}$, the estimated utility is biased $\hat{U}_{\mathcal{A}_{i-1}}(\mathcal{A}_{i}(\mathbf{y}))>U(\mathcal{A}_{i}(\mathbf{y}))$ where $\hat{U}_{\mathcal{A}_{i-1}}$ denotes the utility measured by the previous model and $U$ is the golden utility. The system to optimize for self-preference rather than objective utility, fails to reduce the capability–capacity gap $|\mathbb{D}_{U}-\mathbb{D}_{C}|$. (3) Mode Collapse [167]: Iterative fitting on self-generated data leads to degeneration into a restricted high-frequency sub-distribution, losing generative diversity and associated capability. While $C(\mathcal{A})$ may remain, the capability $U(\mathcal{A})$ can degrade, increasing the misalignment gap $|\mathbb{D}_{U}-\mathbb{D}_{C}|$.
Weak-to-Strong Generalization
While W2SG is relatively practical and requires no human intervention, relying on synthetic signals from weaker $\mathcal{A}_{i-1}$ to supervise stronger $\mathcal{A}_{i}$ can lead to unintended behavior, particularly when there is a significant capability gap $\Delta^{(U)}=|U(\mathcal{A}_{i})-U(\mathcal{A}_{i-1})|$. This stems from two key issues: (1) Imitation of weak supervision [31]. The stronger $\mathcal{A}_{i+1}$ may merely imitate how the weaker $\mathcal{A}_{i}$ behaves, resulting in a near-perfect replication with limited improvement. Also, $\mathcal{A}_{i+1}$ may overfit to the weaker model’s systematic errors or biases, especially when synthetic training signals are flawed. As shown in Fig. 3 and Appendix Fig. 5, we observe that a highly capable AI is more susceptible to such overfitting. Although [41] suggests that larger models tend to agree less with smaller models’ mistakes, such disagreement becomes meaningless in superintelligent tasks where it is infeasible to determine whether $\mathcal{A}_{i}$ or $\mathcal{A}_{i-1}$ is correct. Tab. A.9 illustrates this. When all synthetic labels are flipped, $\mathcal{A}_{i}$ achieves 100% training accuracy but poor test accuracy. This imitation effect occurs when $\Delta^{(U)}$ is large, violating Premise 2. (2) Task salience in the strong AI. W2SG is effective only when the target task $\mathbf{x}$ is internally “salient” to the stronger model $\mathcal{A}_{i+1}$. In tasks where W2SG works well, simple 5-shot prompting tends to perform comparably [31], implying that such capabilities were already accessible and did not require elaborate elicitation. Eliciting non-salient capabilities is more challenging. A small $\delta$ (capability–capacity gap) leads to this problem, where the difficult task is not salient for $\mathcal{A}_{i}$, violating Premise 1.
A.7 Experiment Setup
A.7.1 General
@@
For sandwiching, we use the OpenAI API for the debate experiment. For self-enhancement and weak-to-strong generalization, we fully fine-tune all model variants using 16 NVIDIA A100 GPUs (80GB each), an AMD EPYC 7V12 64-Core Processor CPU, with 1.73TB memory. To reduce memory redundancy during training, we utilize DeepSpeed ZeRO-3 [151] alongside Accelerate [70], using bfloat16 half-precision. We use GPT-2 [146], Phi-3.5 [1], Qwen2.5 [197], and GPT-4 [132] series for our evaluation. For all the model information, please refer to Tab. A.9.
A.7.2 Sandwiching
@@
For the sandwiching experiment, we conduct a debate experiment to evaluate different capabilities in judge-debater and debater-debater setups. In the judge-debater setup, a previous study by Khan et al. [95] demonstrated that the debater is more persuasive when advocating for the correct answer across various debate capabilities. Our goal is to evaluate the effectiveness of debates when the debater and judge have different levels of capability by directly comparing the judge’s accuracy and quality of debate statements.
@@
We use two proprietary models: GPT-4-Turbo and GPT-4o with API access. GPT-4o is known to be a higher-capability model compared to GPT-4-Turbo [133]. We test two different debaters arguing for their assigned answer. Where one debater argues for a correct answer (Helpful), and another debater argues for an incorrect answer (Sneaky). We run 47 randomly chosen GPQA [153] and QuALITY [138] binary QA converted test set for the evaluation with three debate turns. We run each setup four times and calculate the mean values. For QuALITY, we use the debater and judge prompts as described by Khan et al. [95]. For GPQA, we adopt a similar prompt style based on the QuALITY prompts. Detailed templates for GPQA prompts are provided in Tab. LABEL:tab:debater_template, Tab. A.9, Tab. A.9, and Tab. A.9.
@@
A.7.3 Self-Enhancement
For self-enhancement setup, we utilize Learning from AI Feedback (LAIF) training with Direct Preference Optimization (DPO) [148], we adopt a self-training setup to evaluate the performance when there is no supervision from AI systems with greater capabilities. Instead, the supervision comes from models with equal capabilities.